# Optimizing Large Language Models for OpenAPI Code Completion

### v1

@@ V-3 Task-specific fine-tuning can significantly improve the model performance
The 7 billion parameter version of the Code Llama model fine-tuned for OpenAPI completion at the context size of 4 096 tokens following the methods described in Section III-D demonstrates a maximum correctness improvement of 28.6% compared to the original model and further outperforms GitHub Copilot by 55.2%.
V-4 Normal context-level infilling transformation as used in the original Code Llama model can lead to suboptimal performance at smaller context sizes
The case of OpenAPI completion revealed that, when used with documents that are significantly larger than the fine-tuning context size, this technique leads to a bias towards full-context-length completions. This bias can cause the model to generate too long and irrelevant completions when prompted with a context size smaller than the fine-tuning context size. The models trained with this technique still demonstrate the best performance at the context size used during fine-tuning which can be leveraged when the inference context size is known in advance.
@@ V-5 Document splitting technique leads to a more unified performance across context sizes
The document splitting technique proposed in Section III-D can be used to mitigate the problem of the full-context-length infilling bias and lead to more unified performance across the context sizes. Even considering the fact that the models fine-tuned without document splitting outperform the model fine-tuned with this technique at the context size used during fine-tuning, the latter demonstrates the best average performance across the context sizes which is crucial for training foundational or open-source models which are expected to be used with a variety of context sizes.
@@ VI Conclusion
A series of experiments performed in this research demonstrate how modern compact open-source models can be optimized to outperform commercial solutions using the example of a state-of-the-art code generation model from Meta – Code Llama.

### v2

@@ V-3 Task-specific fine-tuning can significantly improve the model performance
The 7-billion-parameter version of the Code Llama model fine-tuned for OpenAPI completion at the context size of 4 096 tokens following the methods described in Section III-D demonstrates a maximum correctness improvement of 28.6% compared to the original model and further outperforms GitHub Copilot by 55.2%.
V-4 Conventional context-level FIM training objective can lead to suboptimal performance at smaller context sizes
The case of OpenAPI completion revealed that, when used with documents that are significantly larger than the fine-tuning context size, this technique leads to a bias towards training-time context size infilling. This bias can cause the model to generate too long and irrelevant completions when prompted with a context size smaller than the fine-tuning context size. The models trained with this technique still demonstrate the best performance at the context size used during fine-tuning which can be leveraged when the inference context size is known in advance.
@@ V-5 Document splitting technique leads to a more unified performance across context sizes
The document splitting technique proposed in Section III-D can be used to mitigate the problem of the training-time context size infilling bias and lead to a more unified performance across the context sizes. Even though the models fine-tuned without document splitting outperform the model fine-tuned with this technique at the context sizes used during fine-tuning, the latter model demonstrates the best average performance across the context sizes which is crucial for training foundational or open-source models that are expected to be used with a variety of context sizes.
@@ VI Conclusion
A series of experiments performed in this research demonstrate how modern compact open-source models can be optimized to outperform commercial solutions. The state-of-the-art code generation model Code Llama from Meta is used here for that.